# mystyle backgroundcolor=white!10

### v1

@@ IV-B Sufficient Conditions for Lower Bounds
Theorem 5
@@
$\begin{cases}\beta+\alpha v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))]&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}\\ v(\bm{x})\leq 1&\forall\bm{x}\in\mathcal{X}_{r}\\ (\alpha-1)v(\bm{x})\leq-\beta&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}\end{cases},$ (13)
@@
$(\alpha-1+\beta)1_{\mathcal{X}_{r}}(\bm{x})+\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\geq\alpha v(\bm{x})+\beta,\forall\bm{x}\in\widehat{\mathcal{X}}.$ (14)
From (14), for $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, we have
@@
The last inequality is obtained via Lemma 2, which states $\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}\geq\mathbb{P}^{\infty}\big{(}\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big{)}$ for $i\leq N$, and the fact that $01-\alpha-\beta<$ 0.
@@
Remark 5
Comparing Theorem 2 and 5, we observe that $\alpha$ cannot be equal to one in Theorem 5. Since if $\alpha=1$, we have $\beta>0$ from $\beta\in(1-\alpha,\infty)$. However, we have $\beta\leq 0$ from $(\alpha-1)v(\bm{x})\leq-\beta,\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}$. This is a contradiction.
V Examples
@@
TABLE I: Computed upper bounds of the probability in Problem 1 in Example 1
($d$ denotes the degree of the polynomial $v(\bm{x})$)
(5) with $\alpha=1.1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.9795 0.9435 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427
(3) with $\alpha=\frac{1}{1.1}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8166 0.1564 0.0650 0.0447 0.0404 0.0398 0.0398 0.0398 0.0398 0.0398
(3) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351
TABLE II: Computed upper bounds of the probability in Problem 2 in Example 1
($d$ denotes the degree of the polynomial $v(\bm{x})$)
(11) with $\alpha=1.1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 0.9868 0.9441 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427
(9) with $\alpha=\frac{1}{1.1}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 0.2530 0.1260 0.0970 0.0906 0.0898 0.0897 0.0897 0.0897 0.0897
(9) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736
Example 1
@@
where $d(\cdot)\colon\mathbb{N}\rightarrow\Theta=[-0.1,0.1]$, $\mathcal{X}=\{\,x\mid h(x)\leq 0\,\}$ with $h(x)=x^{2}-1$, $x_{0}=0.2$, $\mathcal{X}_{r}=\{\,x\mid(x-0.9)^{2}-10^{-4}\leq 0\,\}$, and $N=30$. Besides, we assume that the probability distribution on $\Theta$ is the uniform distribution. The probabilities in Problem 1 and 2 obtained via Monte Carlo methods are around 0.0085 and 0.0128, respectively.
Given the small exact probabilities in Problems 1 and 2, we only estimate their upper bounds in this example. The set $\widetilde{\mathcal{X}}=\widehat{\mathcal{X}}=\{x\mid x^{2}-2\leq 0\}$ is employed in solving (3), (5), (9), and (11). To address these constraints, we encode them into semi-definite programs with sum of squares decomposition techniques for multivariate polynomials. Utilizing polynomials $v(\bm{x})$ of varying degrees, the computed upper bounds are summarized in Tables I and II. Table I illustrates that the constraint (3) with $\alpha=\frac{1}{1.1}$ can yield tighter upper bounds for the probability in Problem 1 compared to (3) with $\alpha=1$ and (5) with $\alpha=1.1$. Similarly, Table II shows that the constraint (9) with $\alpha=\frac{1}{1.1}$ outperforms (9) with $\alpha=1$ and (11) with $\alpha=1.1$ in providing tighter bounds for the probability in Problem 2. Generally, employing higher-degree polynomials for computations facilitates the gain of tighter upper bounds of the probabilities in Problem 1 and 2. However, it is observed that the upper bounds converge rapidly as the polynomial degree increases for this example.
TABLE III: Computed lower and upper bounds of the probability in Problem 1 in Example 2
@@
(5) with $\alpha=1.01$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8798 0.7694 0.7423 0.7302 0.6923 0.6663 0.6130 0.6127 0.5837 0.5830
(5) with $\alpha=1.001$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8169 0.6612 0.6041 0.5845 0.5316 0.4942 0.4131 0.4098 0.3615 0.3605
(3) with $\alpha=\frac{1}{1.01}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1.0000 1.0000 0.9625 0.9235 0.8352 0.7731 0.6342 0.6236 0.5424 0.5397
(3) with $\alpha=\frac{1}{1.001}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8515 0.6877 0.6179 0.5929 0.5381 0.4983 0.4097 0.4027 0.3505 0.3488
(3) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 0.8100 0.6542 0.5880 0.5643 0.5123 0.4745 0.3902 0.3835 0.3345 0.3322
(6) with $\alpha=1.1$ and $\beta=0$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{1}$ $2.5096\times 10^{-14}$ 0.0279 0.0490 0.0502 0.0711 0.0735 0.0990 0.1064 0.1245 0.1289
TABLE IV: Computed lower and upper bounds of the probability in Problem 2 in Example 2
@@
(11) with $\alpha=1.001$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 1 1 1 1 1 1 1 1 1
(11) with $\alpha=1.0001$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 1 1 1 1 1 1 1 1 1
(11) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1 1 1 1 1 1 1 1 1 1
(9) with $\alpha=\frac{1}{1.001}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1.0000 0.9929 0.9827 0.9717 0.9524 0.9380 0.9208 0.9043 0.8932 0.8858
(9) with $\alpha=\frac{1}{1.0001}$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1.0000 0.9506 0.9403 0.9304 0.9117 0.8979 0.8816 0.8658 0.8550 0.8480
(9) with $\alpha=1$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{2}$ 1.0000 0.9460 0.9357 0.9259 0.9073 0.8936 0.8774 0.8617 0.8508 0.8439
(13) with $\alpha=1.06$ and $\beta=0$
d 2 4 6 8 10 12 14 16 18 20
$\epsilon_{1}$ 0.1591 0.2824 0.3453 0.3669 0.4606 0.5218 0.5732 0.5778 0.6119 0.6128
Example 2
@@
where $d(\cdot)\colon\mathbb{N}\rightarrow\Theta=[-1,1]$, $\mathcal{X}=\{\,x\mid h(x)\leq 0\,\}$ with $h(x)=x^{2}-1$, $x_{0}=-0.9$, $\mathcal{X}_{r}=\{\,x\mid x^{2}-0.36\leq 0\,\}$, and $N=50$. Besides, we assume that the probability distribution on $\Theta$ is the uniform distribution. The probabilities in Problem 1 and 2 obtained via Monte Carlo methods are around 0.2321 and 0.7708, respectively.
@@
The set $\widetilde{\mathcal{X}}=\widehat{\mathcal{X}}=\{\,x\mid x^{2}-2.25\leq 0\,\}$ is used in solving (3), (5), (6), (9), (11), and (13). These constraints are addressed via encoding them into semi-definite programs with sum of squares decomposition techniques for multivariate polynomials. Utilizing polynomials $v(\bm{x})$ of varying degrees, the computed lower and upper bounds are summarized in Tables III and IV. It is evident from the results that employing higher-degree polynomials for computations leads to tighter lower and upper bounds of the probabilities in Problems 1 and 2, and the constraints (3) and (5) can complement each other in providing upper bounds of the probability in Problem 1. However, the performance of the constraint (9), with $\alpha=\frac{1}{1.001}$, $\alpha=\frac{1}{1.0001}$, and $\alpha=1$, surpasses that of (11), with $\alpha=1.001$, $\alpha=1.0001$, and $\alpha=1$, in yielding tighter upper bounds of the probability in Problem 2.
@@
In Example 1 and 2, we initially determine upper bounds by assigning pre-defined values to the parameter $\alpha$ in constraints (3), (5), (9) and (11), and lower bounds by assigning pre-defined values to the parameters $\alpha$ and $\beta$ in (6) and (13), and employing convex optimization to solve them. This might yield conservative bounds. However, the automatic optimization of the function $v(\bm{x})$, $\alpha$, and $\beta$ to enhance these bounds constitutes a non-convex problem, which poses a significant challenge. Future work will address this issue, as it is beyond the scope of the current study.
VI Conclusion

### v2

@@ 4.2 Sufficient Conditions for Lower Bounds
The sufficient condition requires a function $v(\bm{x}):\widehat{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying four key properties: (1) its admits a finite upper bound over $\widehat{\mathcal{X}}$, (2) it is less than or equal to one on the target set $\mathcal{X}_{r}$, (3) its expected value at the next step under the dynamics in (1) exceeds its $\alpha$-scaled value by at least $\beta$ for $\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r}$, and (4) it is less than or equal to $\frac{\beta}{\alpha-1}$ over the set $\widehat{\mathcal{X}}\setminus\mathcal{X}$, where $\alpha\in(1,\infty)$ and $\beta\in(1-\alpha,\infty)$. Lower bounds are derived as follows. Lemma 2 equates the probability of system (8) entering $\mathcal{X}_{r}$ at $t=N$ with the cumulative probability of system (1) reaching $\mathcal{X}_{r}$ by time $N$ without leaving $\mathcal{X}$ before hitting the target. Then, we derive a lower bound using system (8). System (8) remains stationary for initial states in $(\widehat{\mathcal{X}}\setminus\mathcal{X})\cup\mathcal{X}_{r}$, implying $\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]=v(\bm{x})$ for $\bm{x}\in(\widehat{\mathcal{X}}\setminus\mathcal{X})\cup\mathcal{X}_{r}$. Given that $v(\bm{x})\leq 1$ on $\mathcal{X}_{r}$ and $v(\bm{x})\leq-\frac{\beta}{\alpha-1}$ over $\widehat{\mathcal{X}}\setminus\mathcal{X}$, we reformulate the constraint over $v(\bm{x})$ using system (8): its expected value at the next step should grow by at least $\beta-(\alpha-1+\beta)1_{\widehat{\mathcal{X}}\setminus\mathcal{X}}(\bm{x})$ relative to its $\alpha$-scaled current value $v(\bm{x})$ for $\bm{x}\in\widehat{\mathcal{X}}$. Recursively applying this reformulated constraint, together with the upper bound on $v$, yields a lower bound.
Theorem 4.
@@
$\begin{cases}\beta+\alpha v(\bm{x})\leq\mathbb{E}^{\infty}[v(\bm{\phi}_{\pi}^{\bm{x}}(1))],&\forall\bm{x}\in\mathcal{X}\setminus\mathcal{X}_{r},\\ v(\bm{x})\leq 1,&\forall\bm{x}\in\mathcal{X}_{r},\\ (\alpha-1)v(\bm{x})\leq-\beta,&\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X},\end{cases}$ (14)
@@
$(\alpha-1+\beta)1_{\mathcal{X}_{r}}(\bm{x})+\mathbb{E}^{\infty}[v(\widehat{\bm{\phi}}_{\pi}^{\bm{x}}(1))]\geq\alpha v(\bm{x})+\beta,\forall\bm{x}\in\widehat{\mathcal{X}}$.
Thus, for $\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}$, we have
@@
The last inequality is obtained via Lemma 2, which states $\mathbb{P}^{\infty}\big(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(N)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big)\geq\mathbb{P}^{\infty}\big(\widehat{\bm{\phi}}^{\bm{x}_{0}}_{\pi}(i)\in\mathcal{X}_{r}\mid\bm{x}_{0}\in\mathcal{X}\setminus\mathcal{X}_{r}\big)$ for $i\leq N$, and the fact that $1-\alpha-\beta<0$.
@@
Remark 5.
Comparing Theorem 2 and 4, we observe that $\alpha$ cannot be equal to one in Theorem 4. Since if $\alpha=1$, we have $\beta>0$ from $\beta\in(1-\alpha,\infty)$. However, we have $\beta\leq 0$ from $(\alpha-1)v(\bm{x})\leq-\beta,\forall\bm{x}\in\widehat{\mathcal{X}}\setminus\mathcal{X}$. This is a contradiction.
5 Examples
@@
Example 1.
@@
where $d(\cdot)\colon\mathbb{N}\rightarrow\Theta=[-0.1,0.1]$, $\mathcal{X}=\{\,x\mid h(x)\leq 0\,\}$ with $h(x)=x^{2}-1$, $\bm{x}_{0}=0.2$, $\mathcal{X}_{r}=\{\,x\mid(x-0.9)^{2}-10^{-4}\leq 0\,\}$, and $N=30$. Besides, we assume that the probability distribution on $\Theta$ is the uniform distribution. The probabilities in Problem 1 and 2 obtained via Monte Carlo methods, which used $10^{4}$ sample paths, are around 0.0085 and 0.0128, respectively.
Given the small exact probabilities in Problems 1 and 2, we only estimate their upper bounds in this example. The set $\widetilde{\mathcal{X}}=\widehat{\mathcal{X}}=\{x\mid x^{2}-2\leq 0\}$ is employed in solving (3), (5), (9), and (11). To address these constraints, we encode them into semi-definite programs with sum of squares decomposition techniques for multivariate polynomials. Utilizing polynomials $v(\bm{x})$ of varying degrees, the computed upper bounds are summarized in Tables 1 and 2. Table 1 illustrates that the constraint (3) with $\alpha=\frac{1}{1.1}$ can yield tighter upper bounds for the probability in Problem 1 compared to (3) with $\alpha=1$ (it also corresponds to the $c$-martingale in [20]) and (5) with $\alpha=1.1$. Similarly, Table 2 shows that the constraint (9) with $\alpha=\frac{1}{1.1}$ outperforms (9) with $\alpha=1$ and (11) with $\alpha=1.1$ in providing tighter bounds for the probability in Problem 2. It is observed that employing higher-degree polynomials for computations facilitates the gain of tighter upper bounds of the probabilities in Problem 1 and 2.
(5) with $\alpha=1.1$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.9795 0.9435 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 (3) with $\alpha=\frac{1}{1.1}$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.8166 0.1564 0.0650 0.0447 0.0404 0.0398 0.0398 0.0398 0.0398 0.0398 (3) with $\alpha=1$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351 0.1351
Table 1: Computed upper bounds of the probability in Problem 1 in Example 1
@@
(11) with $\alpha=1.1$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.9943 0.9553 0.9428 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 0.9427 (9) with $\alpha=\frac{1}{1.1}$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 1 0.2530 0.1260 0.0970 0.0906 0.0898 0.0897 0.0897 0.0897 0.0897 (9) with $\alpha=1$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736 0.1736
Table 2: Computed upper bounds of the probability in Problem 2 in Example 1
@@
(5) with $\alpha=1.01$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.8798 0.7694 0.7423 0.7302 0.6923 0.6663 0.6130 0.6127 0.5837 0.5830 (5) with $\alpha=1.001$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.8169 0.6612 0.6041 0.5845 0.5316 0.4942 0.4131 0.4098 0.3615 0.3605 (3) with $\alpha=\frac{1}{1.01}$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 1.0000 1.0000 0.9625 0.9235 0.8352 0.7731 0.6342 0.6236 0.5424 0.5397 (3) with $\alpha=\frac{1}{1.001}$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.8515 0.6877 0.6179 0.5929 0.5381 0.4983 0.4097 0.4027 0.3505 0.3488 (3) with $\alpha=1$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 0.8100 0.6542 0.5880 0.5643 0.5123 0.4745 0.3902 0.3835 0.3345 0.3322 (6) with $\alpha=1.1$ and $\beta=0$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{1}$ $2.5096\times 10^{-14}$ 0.0279 0.0490 0.0502 0.0711 0.0735 0.0990 0.1064 0.1245 0.1289
Table 3: Computed lower and upper bounds of the probability in Problem 1 in Example 2
($d$ denotes the degree of the polynomial $v(\bm{x})$)
(11) with $\alpha=1.001$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 1 0.9485 0.9388 0.9264 0.9117 0.8988 0.8832 0.8682 0.8579 0.8513 (11) with $\alpha=1.0001$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 1 0.9463 0.9360 0.9263 0.9077 0.8941 0.8780 0.8623 0.8515 0.8446 (11) with $\alpha=1$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 1.0000 0.9460 0.9357 0.9259 0.9073 0.8936 0.8774 0.8617 0.8508 0.8439 (9) with $\alpha=\frac{1}{1.001}$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 1.0000 0.9929 0.9827 0.9717 0.9524 0.9380 0.9208 0.9043 0.8932 0.8858 (9) with $\alpha=\frac{1}{1.0001}$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 1.0000 0.9506 0.9403 0.9304 0.9117 0.8979 0.8816 0.8658 0.8550 0.8480 (9) with $\alpha=1$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{2}$ 1.0000 0.9460 0.9357 0.9259 0.9073 0.8936 0.8774 0.8617 0.8508 0.8439 (14) with $\alpha=1.06$ and $\beta=0$ d 2 4 6 8 10 12 14 16 18 20 $\epsilon_{1}$ 0.1591 0.2824 0.3453 0.3669 0.4606 0.5218 0.5732 0.5778 0.6119 0.6128
Table 4: Computed lower and upper bounds of the probability in Problem 2 in Example 2
($d$ denotes the degree of the polynomial $v(\bm{x})$)
Example 2.
@@
where $d(\cdot)\colon\mathbb{N}\rightarrow\Theta=[-1,1]$, $\mathcal{X}=\{\,x\mid h(x)\leq 0\,\}$ with $h(x)=x^{2}-1$, $\bm{x}_{0}=-0.9$, $\mathcal{X}_{r}=\{\,x\mid x^{2}-0.36\leq 0\,\}$, and $N=50$. Besides, we assume that the probability distribution on $\Theta$ is the uniform distribution. The probabilities in Problem 1 and 2 obtained via Monte Carlo methods, which used $10^{4}$ sample paths, are around 0.2321 and 0.7708, respectively.
@@
The set $\widetilde{\mathcal{X}}=\widehat{\mathcal{X}}=\{\,x\mid x^{2}-2.25\leq 0\,\}$ is used in solving (3), (5), (6), (9), (11), and (14). These constraints are addressed via encoding them into semi-definite programs with sum of squares decomposition techniques for multivariate polynomials. Utilizing polynomials $v(\bm{x})$ of varying degrees, the computed lower and upper bounds are summarized in Tables 3 and 4. It is evident from the results that employing higher-degree polynomials for computations leads to tighter lower and upper bounds of the probabilities in Problems 1 and 2, and the constraints (3) and (5) can complement each other in providing upper bounds of the probability in Problem 1. However, the performance of the constraint (11), with $\alpha=\frac{1}{1.001}$, $\alpha=\frac{1}{1.0001}$, marginally surpasses that of (9), with $\alpha=1.001$, $\alpha=1.0001$, in yielding tighter upper bounds for the probability in Problem 2.
@@
In Example 1 and 2, we initially determine upper bounds by assigning pre-defined values to the parameter $\alpha$ in constraints (3), (5), (9) and (11), and lower bounds by assigning pre-defined values to the parameters $\alpha$ and $\beta$ in (6) and (14), and employing convex optimization to solve them. This might yield conservative bounds. However, the automatic optimization of the function $v(\bm{x})$, $\alpha$, and $\beta$ to enhance these bounds constitutes a non-convex problem, which poses a significant challenge. Future work will address this issue, as it is beyond the scope of the current study. Furthermore, this work does not concentrate on the design of efficient algorithms to solve the associated constraints. Instead, it leverages existing semi-definite programming tools to tackle these issues. Future work will address this gap by proposing efficient algorithms to effectively address these constraints.
6 Conclusion